Ground state energy of a particle in a one dimensional infinite potential box is 2 ev then energy of first excited state is how much
1
Expert's answer
2021-09-30T10:19:46-0400

Energy of a particle in the ground is "E_1=2.0 e.V"

The energy of a particle in the first excited state is

"E_2=2^2\times E_1\\

=2^2\times(2.0\; eV)\\

=8 \; eV"